Suppose X1 and X2 are iid Poisson(θ) random variables and let T = X1 + 2X2.

(a) Find the conditional distribution of (X1,X2) given T = 7.

(b) For θ = 1 and θ = 2, respectively, calculate all probabilities in the above conditional distribution and present the two conditional distributions numerically.
Answer #1

if x1+ 2x2 = 7 P(Xi = xi , X2 = (7-x1 )/2) P(X1-2X2-7) 1) + P(Xi = 7)P(X2 since Xi and X2 are independent T-I105%0 when r1 1 12 92 120 6 12 т 120-5040 5040 when r15 wien x1 = 7 = 0 otherwise (b) For θ-1 P(X1 = x1|T = 7) = 0.6447 when x